Consumer spends $360 per week on two goods X and Y. PX=$3 and PY=$2.Hisutility

functionisU=2X2Y.

What quantities of X and Y does he buy each week in equilibrium?.



1
Expert's answer
2021-03-31T11:46:51-0400

we have: I=$360, U=2X2Y, PX=$3, PY=$2

thus "U_{xx}={ d\over dx}({ dU\over dx})=4Y, U_{yy}={ d\over dy}({ dU\over dy})=0"

"U_{xy}={ d\over dx}({ dU\over dy})=4X={ d\over dy}({ dU\over dx})=U_{yx}"

"U_{xx}={ d\over dx}({ dU\over dx})=4Y, U_{yy}={ d\over dy}({ dU\over dy})=0"

At equilibrium

"{ MU_x\over MU_y}={ 4XY\over 2X^2}={ P_x\over P_y}\implies2YP_y=XP_x"

"YP_y+XP_x=I"

"3YP_y=I"

"Y={ I\over 3P_x}=60"

"YP_y={ 1\over 2}XP_x\implies{ 1\over 2}XP_x+XP_x=I"

"\implies X={ 2I\over 3P_x}=80"